# citations.bib

### v1

@@ A.1 Notation
[…]
@@ A.2 Proofs from Randomized Numerical Linear Algebra
In this section, we provide bounds for approximate matrix-multiplication and sketched regression. Our proofs follow standard techniques in randomized numerical linear algebra [woodruff2014sketching, Martinsson2020, Tropp2015], and are included to illustrate core concepts which may provide a useful starting point for proving theoretical guarantees for more complicated sketching distributions for Shapley value estimation. For simplicity, we analyze the simpler case that $\bm{S}$ has independent rows; i.e. that
@@
where $I_{1},\ldots,I_{m}$ are iid copies of a random variable $I$ for which $\mathbb{P}[I=k]=p_{k}$, $k\in[r]$ for some fixed $r\in\mathbb{N}$.
Proofs for sampling without replacement, which follow a similar strategy, are described in Section A.5.
A.2.1 Approximate Matrix Multiplication
We begin with a simple bound on approximate matrix multiplication. This bound immediately gives provable guarantees for the Shapley estimator $\bm{\phi}^{\textup{M}}$ defined in Section 2.1.
Theorem A.1 (Matrix multiplication).
@@ A.2.1 Approximate Matrix Multiplication
Let $I_{1},\dotsc,I_{m}$ denote $m$ iid random variables that sample indices from $[r]$ according to probability $\mathcal{P}$. Let $\bm{u}_{1},\dotsc,\bm{u}_{r}$ be the columns of $\bm{U}^{\mathsf{T}}$, and define $\bm{X}_{j}=\bm{u}_{I_{j}}z_{I_{j}}/p_{I_{j}}$ for $j\in[m]$. Then, $\bm{X}_{j}$ are iid $d$-dimensional random vectors. It can be verified that $\mathbb{E}[\bm{X}_{j}]=\bm{U}^{\mathsf{T}}\bm{z}$ for all $j\in[m]$. Next, we calculate the variance of the random vector $\bm{X}_{j}$ for $j\in[m]$. Using Eq. 2.3, observe that
@@ A.2.2 Subspace Embedding
Before we prove a bound for the sketched regression $\bm{\phi}^{\textup{R}}$ from Section 2.1, we prove a subspace embedding guarantee.
We begin by recalling the following well-known matrix concentration inequality; see e.g., [Tropp2015, Theorem 6.6.1].
@@
Writing $\sum_{i=1}^{m}\mathbb{E}\left[\bm{X}_{i}^{2}\right]=m\sigma^{2}$, we have
@@
The result then follows from A.2. ∎
@@ A.2.3 Sketched Regression
Together, Theorems A.1 and A.3 give a bound on sketched regression.
@@
Since $\bm{y}^{*}$ is the solution of $\min_{\bm{y}}\left\lVert\bm{U}\bm{y}-\bm{b}\right\rVert$, $\bm{U}\bm{y}^{*}-\bm{b}$ lies in the orthogonal complement of the range of $\bm{U}$, and therefore, $\bm{U}^{\mathsf{T}}(\bm{U}\bm{y}^{*}-\bm{b})=0$. Then, taking $\bm{z}=\bm{U}\bm{y}^{*}-\bm{b}=(\bm{I}-\bm{U}\bm{U}^{\mathsf{T}})\bm{b}$ in Theorem A.1, we can infer that using
@@
we have with probability exceeding $21-\delta/$ 2,
@@
$\left\lVert\sum_{i=1}^{r}\frac{\left\lVert\bm{u}_{i}\right\rVert^{2}}{p_{i}}\bm{u}_{i}\bm{u}_{i}^{\mathsf{T}}-(\bm{U}^{\mathsf{T}}\bm{U})^{2}\right\rVert\leq(\eta-1)\left\lVert\bm{U}^{\mathsf{T}}\bm{U}\right\rVert\leq\eta.$ (A.14)
Therefore, by Theorem A.3, if
@@
then, with probability exceeding $21-\delta/$ 2,
@@
For the remainder of the proof, we condition on Eqs. A.12 and A.17, which, by a union bound, simultaneously occur with probability exceeding $1-\delta$.
Using the triangle inequality, submultiplicitivty, and Eq. A.17,
$\left\lVert\bm{y}^{*}-\widehat{\bm{y}}\right\rVert$ $=\left\lVert(\bm{U}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{S}\bm{U})(\bm{y}^{*}-\widehat{\bm{y}})+(\bm{I}-(\bm{U}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{S}\bm{U}))(\bm{y}^{*}-\widehat{\bm{y}})\right\rVert$ (A.18)
$\leq\left\lVert(\bm{U}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{S}\bm{U})(\bm{y}^{*}-\widehat{\bm{y}})\right\rVert+\left\lVert(\bm{I}-(\bm{U}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{S}\bm{U}))(\bm{y}^{*}-\widehat{\bm{y}})\right\rVert$ (A.19)
@@
$\left\lVert\bm{y}^{*}-\widehat{\bm{y}}\right\rVert\leq 2\left\lVert(\bm{U}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{S}\bm{U})(\bm{y}^{*}-\widehat{\bm{y}})\right\rVert.$ (A.22)
Next, by the optimality of $\widehat{\bm{y}}$ we have that $(\bm{S}\bm{U})^{\mathsf{T}}(\bm{S}\bm{U}\bm{\widehat{y}}-\bm{S}\bm{b})=\bm{0}$ and hence that $(\bm{U}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{S}\bm{U})\widehat{\bm{y}}=\bm{U}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{S}\bm{b}$. Therefore, by Eq. A.12,
@@ A.3 Proofs for Shapley Value Estimators
As noted in [musco2025provably, Lemma 3.3], the matrix $\bm{Z}^{\prime}$ nearly has orthonormal columns.
@@
Let $c_{d}={((d-1)H_{d-2}-(d-2))}/{d}$, where $H_{d}$ is the $d^{th}$ harmonic number. Then,
@@
$=\frac{1}{d}\sum_{|S|=2}^{d}\frac{|S|-1}{d-|S|}=\frac{1}{d}\sum_{|S|=2}^{d}\frac{d-1}{d-|S|}-1$ (A.31)
@@
Next, we describe the conversion from the constrained problem Eq. 1.2 to an unconstrained problem. Our approach is closely related to [musco2025provably, Lemma 3.1]. However, as noted in Section 2, our approach allows arbitrary $\lambda$ (where as [musco2025provably] only allows $\lambda=\alpha$). In addition, on a more technical note, we state our results in terms of the argmin of a regression problem involving a full-rank matrix $\bm{U}$. The result of [musco2025provably] is stated in terms of the argmin of a regression problem involving rank-deficient matrix, which is not uniquely defined. As such, their result implicitly assumes that the argmin returns one particular solution (the minimum norm solution); see Section B.3.
Proof of Theorem 2.1.
Next, using that $\bm{Q}^{\mathsf{T}}\bm{1}=\bm{0}$ and that $\bm{Q}^{\mathsf{T}}\bm{Q}=\bm{I}$,
@@
Therefore $(\bm{U}^{\mathsf{T}}\bm{U})^{-1}\bm{U}^{\mathsf{T}}=\bm{U}^{\mathsf{T}}$ and so $\bm{Q}\operatornamewithlimits{argmin}_{\bm{x}\in\mathbb{R}^{d-1}}\left\lVert\bm{U}\bm{x}-\bm{b}_{\lambda}\right\rVert^{2}+\alpha\bm{1}=\bm{Q}\bm{U}^{\mathsf{T}}\bm{b}_{\lambda}+\alpha\bm{1}$.
It remains to show these formulations are equivalent to Eq. 1.2. Since $\bm{Q}^{\mathsf{T}}\bm{1}=\bm{0}$ and $\bm{1}^{\mathsf{T}}\bm{1}=d$, observe that
@@
$\bm{\phi}^{*}$ $=\operatornamewithlimits{argmin}_{\begin{subarray}{c}\bm{\phi}\in\mathbb{R}^{d}\\ \bm{1}^{\mathsf{T}}\bm{\phi}=v([d])-v(\emptyset)\end{subarray}}\left\lVert\bm{Z}^{\prime}\bm{\phi}-\bm{b}\right\rVert$ (A.36)
@@
$\operatornamewithlimits{argmin}_{\bm{x}\in\mathbb{R}^{d-1}}\lVert\bm{U}\bm{x}-\bm{b}_{\lambda}\rVert=\operatornamewithlimits{argmin}_{\bm{x}\in\mathbb{R}^{d-1}}\lVert\bm{U}\bm{x}-(\bm{b}-\lambda\bm{Z}^{\prime}\bm{1})\rVert=\operatornamewithlimits{argmin}_{\bm{x}\in\mathbb{R}^{d-1}}\lVert\bm{U}\bm{x}-\bm{b}\rVert.$ (A.40)
@@
Finally, we use Theorem 2.1 and the bounds from Section A.2 to prove our main approximation guarantee.
Proof of Theorem 2.2.
We analyze the estimators individually. Recall from Theorem 2.1 that
@@
$\bm{\phi}_{\lambda}^{\textup{R}}=\bm{Q}\operatornamewithlimits{argmin}_{\bm{x}\in\mathbb{R}^{d-1}}\left\lVert\bm{S}(\bm{U}\bm{x}-\bm{b}_{\lambda})\right\rVert+\alpha\bm{1}=\bm{Q}(\bm{S}\bm{U})^{+}\bm{S}\bm{b}_{\lambda}+\alpha\bm{1}.$
@@
Matrix Multiplication: By definition,

### v2

@@ A.1 Notation
Unless mentioned otherwise, $\bm{e}_{1},\dotsc,\bm{e}_{q}$ are the standard basis vectors for $\mathbb{R}^{q}$. $\bm{0}$ and $\bm{1}$ are the vectors of all zeros and ones, respectively, while $\bm{I}$ is the identity matrix. $\lVert\cdot\rVert$ denotes the Euclidean norm for vectors, while the spectral norm for matrices. $\lVert\cdot\rVert_{\mathsf{F}}$ denotes the Frobenius norm. Given a matrix $\bm{A}$, $\bm{A}^{+}$ denotes its Moore-Penrose pseudoinverse. Finally, given a matrix $\bm{U}$ with orthonormal columns, we write $\bm{P}_{\bm{U}}=\bm{I}-\bm{U}\bm{U}^{\mathsf{T}}$.
@@ A.2 Proofs from Randomized Numerical Linear Algebra
In this section, we provide bounds for approximate matrix-vector multiplication and sketched regression (least squares) as defined in Section˜2.1. Our proofs follow standard techniques in randomized numerical linear algebra [Woo+14, MT20, Tro15], and are included to illustrate core concepts which may provide a useful starting point for proving theoretical guarantees for more complicated sketching distributions for Shapley value estimation. For simplicity, we analyze the simpler case that $\bm{S}$ has independent rows; i.e. that
@@
where $I_{1},\ldots,I_{m}$ are iid copies of a random variable $I$ for which $\mathbb{P}[I=k]=p_{k}$, $k\in[r]$ for some fixed $r\in\mathbb{N}$. Note that $\mathbb{E}[\bm{S}^{\mathsf{T}}\bm{S}]=\bm{I}$.
Proof Sketch: Prior to diving into the technical details, we present a high-level overview of the strategy used in deriving sample complexity bounds for matrix-vector multiplication and regression estimators.
Given a matrix $\bm{U}\in\mathbb{R}^{r\times q}$ satisfying $\bm{U}^{\mathsf{T}}\bm{U}=\bm{I}$, fixed vectors $\bm{z},\bm{b}\in\mathbb{R}^{r}$, our goal is to estimate $\bm{U}^{\mathsf{T}}\bm{z}$ and $\bm{y}^{*}=\operatornamewithlimits{argmin}_{\bm{y}}\left\lVert\bm{U}\bm{y}-\bm{b}\right\rVert^{2}$.
  1. 1.
    Approximate Matrix-Vector Multiplication: Observe that $\bm{U}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{S}\bm{z}$ is an unbiased estimator of $\bm{U}^{\mathsf{T}}\bm{z}$. By computing the variance of this estimator and using Markov’s inequality, we obtain bounds on the sample complexity of estimating $\bm{U}^{\mathsf{T}}\bm{z}$ to a given error (in $\ell_{2}$ norm) and confidence level (see Theorem˜A.1). Note, in particular, that the term $\gamma(\bm{z})$ (see Eq.˜2.3) appearing in the sample complexity is related to the variance of the estimator.
  2. 2.
    Sketched Regression: The sketched regression estimator is given by $\widehat{\bm{y}}=\operatornamewithlimits{argmin}_{\bm{y}}\left\lVert\bm{S}(\bm{U}\bm{y}-\bm{b})\right\rVert^{2}$. To derive sample complexity bounds for estimating $\bm{y}^{*}$ using $\widehat{\bm{y}}$, we use two main observations. (I) Since $\bm{U}^{\mathsf{T}}\bm{S}^{T}\bm{S}\bm{U}$ is an unbiased estimator of $\bm{I}$, we can compute the number of samples (using matrix Bernstein’s inequality, ˜A.3) to ensure that $\left\lVert\bm{U}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{S}\bm{U}-\bm{I}\right\rVert\leq 1/2$ holds with sufficiently high probability (see Theorem˜A.4). Straightforward algebra then gives $\left\lVert\hat{\bm{y}}-\bm{y}^{*}\right\rVert\leq 2\left\lVert\bm{U}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{S}\bm{U}(\widehat{\bm{y}}-\bm{y}^{*})\right\rVert$ with high probability. (II) Since $\bm{S}(\bm{U}\widehat{\bm{y}}-\bm{b})$ lies in the orthogonal complement of $\bm{S}\bm{U}$, we have $\bm{U}^{\mathsf{T}}\bm{S}^{\mathsf{T}}(\bm{S}\bm{U}\widehat{\bm{y}}-\bm{S}\bm{b})=0$. For a similar reason, we also have $\bm{U}^{\mathsf{T}}(\bm{U}\bm{y}^{*}-\bm{b})=0$. It follows that $\bm{U}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{S}\bm{U}(\bm{y}^{*}-\widehat{\bm{y}})=\bm{U}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{S}(\bm{U}\bm{y}^{*}-\bm{b})$, which is just a sketched matrix-vector multiplication estimator for $\bm{U}^{\mathsf{T}}(\bm{U}\bm{y}^{*}-\bm{b})=0$. Consequently, we can use Theorem˜A.1 to compute the sample complexity for bounding the error $\left\lVert\bm{U}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{S}\bm{U}(\widehat{\bm{y}}-\bm{y}^{*})\right\rVert$ with high probability. In particular, since $\bm{U}^{\mathsf{T}}(\bm{U}\bm{y}^{*}-\bm{b})=(\bm{I}-\bm{U}\bm{U}^{\mathsf{T}})\bm{b}=\bm{P}_{\bm{U}}\bm{b}$, this explains why we have $\gamma(\bm{P}_{\bm{U}}\bm{b})$ instead of $\gamma(\bm{b})$ in the sample complexity for the sketched regression estimator in Theorem˜A.5.
Proofs for sampling without replacement, which follow a similar strategy, are described in Section˜A.5.
A.2.1 Approximate Matrix-Vector Multiplication
We begin with a simple bound on approximate matrix-vector multiplication. This bound immediately gives provable guarantees for the Shapley estimator $\bm{\phi}^{\textup{M}}$ defined in Section˜2.1.
Theorem A.1 (Matrix-Vector multiplication).
@@ A.2.1 Approximate Matrix-Vector Multiplication
Let $I_{1},\dotsc,I_{m}$ denote $m$ iid random variables that sample indices from $[r]$ according to probability $\mathcal{P}$. Let $\bm{u}_{1},\dotsc,\bm{u}_{r}$ be the columns of $\bm{U}^{\mathsf{T}}$, and define $\bm{X}_{j}=\bm{u}_{I_{j}}z_{I_{j}}/p_{I_{j}}$ for $j\in[m]$. Then, $\bm{X}_{j}$ are iid $d$-dimensional random vectors. It can be verified that $\mathbb{E}[\bm{X}_{j}]=\bm{U}^{\mathsf{T}}\bm{z}$ for all $j\in[m]$. Next, we calculate the variance of the random vector $\bm{X}_{j}$ for $j\in[m]$. Using Eq.˜2.3, observe that
@@
Remark A.2.
When using Theorem˜A.1 to compute the sample complexity bound in the subsequent proofs in Appendix˜A, we use
$m=O\left(\frac{\gamma(\bm{z})}{\delta\varepsilon^{2}}\right)$ (A.5)
samples. This is, in general, an upper bound on the sample complexity required for approximating the matrix-vector product, and can be tightened by including the term $\left\lVert\bm{U}^{\mathsf{T}}\bm{z}\right\rVert^{2}$.
@@ A.2.2 Subspace Embedding
Before we prove a bound for the sketched regression $\bm{\phi}^{\textup{R}}$ from Section˜2.1, we prove a subspace embedding guarantee.
We begin by recalling the following well-known matrix concentration inequality; see e.g., [Tro15, Theorem 6.6.1].
@@
Writing $\left\lVert\sum_{i=1}^{m}\mathbb{E}\left[\bm{X}_{i}^{2}\right]\right\rVert=m\sigma^{2}$, we have
@@
The result then follows from ˜A.3. ∎
@@ A.2.3 Sketched Regression
Together, Theorems˜A.1 and A.4 give a bound on sketched regression.
@@
Since $\bm{y}^{*}$ is the solution of $\min_{\bm{y}}\left\lVert\bm{U}\bm{y}-\bm{b}\right\rVert^{2}$, $\bm{U}\bm{y}^{*}-\bm{b}$ lies in the orthogonal complement of the range of $\bm{U}$, and therefore, $\bm{U}^{\mathsf{T}}(\bm{U}\bm{y}^{*}-\bm{b})=0$. Then, taking $\bm{z}=\bm{U}\bm{y}^{*}-\bm{b}=(\bm{I}-\bm{U}\bm{U}^{\mathsf{T}})\bm{b}$ in Theorem˜A.1, we can infer that using
@@
we have with probability exceeding $1-\delta/2$,
@@
$\left\lVert\sum_{i=1}^{r}\frac{\left\lVert\bm{u}_{i}\right\rVert^{2}}{p_{i}}\bm{u}_{i}\bm{u}_{i}^{\mathsf{T}}-(\bm{U}^{\mathsf{T}}\bm{U})^{2}\right\rVert\leq(\eta-1)\left\lVert\bm{U}^{\mathsf{T}}\bm{U}\right\rVert=\eta-1\leq\eta.$ (A.15)
Therefore, by Theorem˜A.4, if
@@
then, with probability exceeding $1-\delta/2$,
@@
For the remainder of the proof, we condition on Eqs.˜A.13 and A.18, which, by a union bound, simultaneously occur with probability exceeding $1-\delta$.
Using the triangle inequality, submultiplicitivty, and Eq.˜A.18,
[…]
@@
[…]
Next, by the optimality of $\widehat{\bm{y}}$ we have that $(\bm{S}\bm{U})^{\mathsf{T}}(\bm{S}\bm{U}\bm{\widehat{y}}-\bm{S}\bm{b})=\bm{0}$ and hence that $(\bm{U}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{S}\bm{U})\widehat{\bm{y}}=\bm{U}^{\mathsf{T}}\bm{S}^{\mathsf{T}}\bm{S}\bm{b}$. Therefore, by Eq.˜A.13,
@@ A.3 Proofs for Shapley Value Estimators
As noted in [MW25, Lemma 3.3], the matrix $\bm{Z}^{\prime}$ nearly has orthonormal columns.
@@
Let $c_{d}={((d-1)H_{d-2}-(d-2))}/{d}$, where $H_{d}=\sum_{i=1}^{d}(1/i)$ is the $d^{th}$ harmonic number. Then,
@@
$=\frac{1}{d}\sum_{|S|=2}^{d-1}\frac{|S|-1}{d-|S|}$ (A.32)
$=\frac{1}{d}\sum_{|S|=2}^{d-1}\left(\frac{d-1}{d-|S|}-1\right)$ (A.33)
@@
Next, we describe the conversion from the constrained problem Eq.˜1.2 to an unconstrained problem. Our approach is closely related to [MW25, Lemma 3.1]. However, as noted in Section˜2, our approach allows arbitrary $\lambda$ (where as [MW25] only allows $\lambda=\alpha$). In addition, on a more technical note, we state our results in terms of the argmin of a regression problem involving a full-rank matrix $\bm{U}$. The result of [MW25] is stated in terms of the argmin of a regression problem involving rank-deficient matrix, which is not uniquely defined. As such, their result implicitly assumes that the argmin returns one particular solution (the minimum norm solution); see Section˜B.3.
Proof of Theorem˜2.1.
Using $\bm{Q}^{\mathsf{T}}\bm{1}=\bm{0}$, $\bm{Q}^{\mathsf{T}}\bm{Q}=\bm{I}$, and Eq.˜A.25,
@@
Therefore, $(\bm{U}^{\mathsf{T}}\bm{U})^{-1}\bm{U}^{\mathsf{T}}=\bm{U}^{\mathsf{T}}$ and so $\bm{Q}\operatornamewithlimits{argmin}_{\bm{x}\in\mathbb{R}^{d-1}}\left\lVert\bm{U}\bm{x}-\bm{b}_{\lambda}\right\rVert^{2}+\alpha\bm{1}=\bm{Q}\bm{U}^{\mathsf{T}}\bm{b}_{\lambda}+\alpha\bm{1}$.
It remains to show these formulations are equivalent to Eq.˜1.2. Since $\bm{Q}^{\mathsf{T}}\bm{1}=\bm{0}$ and $\bm{1}^{\mathsf{T}}\bm{1}=d$, observe that
@@
$\bm{\phi}^{*}$ $=\operatornamewithlimits{argmin}_{\begin{subarray}{c}\bm{\phi}\in\mathbb{R}^{d}\\ \bm{1}^{\mathsf{T}}\bm{\phi}=v([d])-v(\emptyset)\end{subarray}}\left\lVert\bm{Z}^{\prime}\bm{\phi}-\bm{b}\right\rVert^{2}$ (A.38)
@@
$\operatornamewithlimits{argmin}_{\bm{x}\in\mathbb{R}^{d-1}}\lVert\bm{U}\bm{x}-\bm{b}_{\lambda}\rVert^{2}=\operatornamewithlimits{argmin}_{\bm{x}\in\mathbb{R}^{d-1}}\lVert\bm{U}\bm{x}-(\bm{b}-\lambda\bm{Z}^{\prime}\bm{1})\rVert^{2}=\operatornamewithlimits{argmin}_{\bm{x}\in\mathbb{R}^{d-1}}\lVert\bm{U}\bm{x}-\bm{b}\rVert^{2}.$ (A.42)
@@
Finally, we use Theorem˜2.1 and the bounds from Section˜A.2 to prove our main approximation guarantee.
Proof of Theorem˜2.2.
We analyze the estimators individually. Recall from Theorem˜2.1 that
@@
$\bm{\phi}_{\lambda}^{\textup{R}}=\bm{Q}\operatornamewithlimits{argmin}_{\bm{x}\in\mathbb{R}^{d-1}}\left\lVert\bm{S}(\bm{U}\bm{x}-\bm{b}_{\lambda})\right\rVert^{2}+\alpha\bm{1}=\bm{Q}(\bm{S}\bm{U})^{+}\bm{S}\bm{b}_{\lambda}+\alpha\bm{1}.$
@@
Matrix-Vector Multiplication: By definition,